### v1

@@ 5.5 Evaluation
Figure 8: Examples of datasets, EgoPER [lee2024error] and CaptainCook4D [peddi2024captaincook4d] focus on kitchen cooking scenarios, EgoOops [haneji2024egooops] designs complex scenarios such as chemical experiments and building blocks, and EgoproceL [bansal2022my] focuses on pc assembly and disassembly tasks.
@@ 6 Experiments
Table 7: Experimental results in egocentric procedural error detection, the part highlighted in purple is the result of our supplementary experiments, bold text represents the best performance on this metric in each part.
Method Years Precision EDA AUC F1 Recall Acc Dataset
EgoPED [lee2024error] CVPR 2024 56.5 69.8 54.9 - - 69.8 CaptainCook4D
AMNAR [huang2025modeling] CVPR 2025 66.8 72.3 60.2 - - 72.3
Just Ask [yang2021just] ICCV 2021 35.1 35.1 50.0 51.2 - 35.1
Vinci [huang2024vinci] arXiv 2024 35.6 49.7 48.5 - - 49.7
Video-LLaVA [lin2023video] arXiv 2024 40.1 60.8 52.2 29.6 23.4 60.8
Video-LLaMA2 [cheng2024videollama] arXiv 2024 36.5 58.4 50.1 29.8 25.2 58.4
LLaVA-OneVision [li2024llava] arXiv 2024 39.8 63.0 53.1 17.0 10.8 63.0
EgoGPT [yang2025egolife] CVPR 2025 36.0 46.5 49.6 47.0 67.5 46.5
Vinci [huang2024vinci] arXiv 2024 19.4 26.9 47.3 30.8 74.3 26.9 EgoOops
Video-LLaVA [lin2023video] arXiv 2024 14.6 60.0 44.6 15.8 17.1 60.0
Video-LLaMA2 [cheng2024videollama] arXiv 2024 15.4 50.0 42.4 20.0 28.6 50.0
LLaVA-OneVision [li2024llava] arXiv 2024 26.1 64.4 54.1 29.6 34.3 64.4
EgoGPT [yang2025egolife] CVPR 2025 34.8 73.8 55.1 27.6 22.9 73.8
Vinci [huang2024vinci] arXiv 2024 12.0 49.3 47.4 19.3 49.2 49.3 EgoPER
Video-LLaVA [lin2023video] arXiv 2024 13.5 77.5 50.7 14.3 15.3 77.5
Video-LLaMA2 [cheng2024videollama] arXiv 2024 13.6 75.2 51.6 15.8 18.9 75.2
LLaVA-OneVision [li2024llava] arXiv 2024 14.0 60.1 54.5 21.3 43.7 60.1
EgoGPT [yang2025egolife] CVPR 2025 14.4 69.6 51.3 19.4 29.8 69.6
@@ 6.1 Datasets and Implementation Details
Datasets. To ensure broad scenario coverage, three egocentric datasets named CaptainCook4D [peddi2024captaincook4d], EgoPER [lee2024error], and EgoOops [haneji2024egooops] were employed for the egocentric procedural error detection task: CaptainCook4D and EgoPER address diverse kitchen activities, while EgoOops targets niche contexts such as chemical experiments and block-building. As these datasets lack predefined test sets, 20% of CaptainCook4D and 30% of EgoOops samples were randomly designated as test sets. For the procedural learning task, EgoProceL [bansal2022my] was used, with PC assembly and disassembly tasks selected to assess the AI assistant’s competence in extracting critical procedural steps.
Implementation Details. Both video question answering models and AI assistants are evaluated using question answering tasks. For egocentric procedural error detection, each model receives 4 to 5 randomly selected questions, with prompts tailored to highlight the error detection objective and specify potential error categories, thereby maximizing detection performance without fine-tuning. Following the protocol described in [huang2025modeling], keywords from model outputs are parsed to generate confidence scores, which are then used to calculate Precision, EDA, AUC, F1, Recall, and Accuracy. This comprehensive metric set ensures a robust assessment of error detection. Consistent prompt formats and confidence rules are applied across EgoPER and EgoOops, with minor adjustments for CaptainCook4D to accommodate model-specific attributes. For egocentric procedural learning, five models were systematically evaluated on the PC assembly and PC disassembly tasks within EgoproceL, adhering to the experimental protocol established by [mahmood2025procedure]. Performance was quantitatively assessed using the F1 score and Intersection over Union (IoU) metrics, and the results were benchmarked against existing procedural learning methodologies. To address relevant task requirements, tailored prompts were introduced. Given that the models lacked fine-tuning for key step learning and acknowledging the inherent challenge in uniformly evaluating semantically similar key step sequences, the egocentric procedural learning task was suitably simplified. To mitigate excessive deviation between model predictions and expected outputs, a predefined set of candidate key steps was embedded in the prompts, thereby enabling the models to select the sequence of steps most closely matching each video segment. In accordance with the specifications of [mahmood2025procedure], and to enhance the objectivity of evaluating procedural learning abilities, the Hungarian matching algorithm was employed to analyze the result.

### v2

@@ 5.5 Evaluation
Figure 7: (a) shows the different methods in real-time and streaming video understanding: VideoLLM-online [chen2024videollm], Flash-VStream [zhang2024flash], Dispider [qian2025dispider], StreamBridge [wang2025streambridge], StreamChat [xiongstreaming], MMDuet [wang2024videollm], and (b) shows that in proactive interaction in procedural contexts: YETI [bandyopadhyay2025yeti], PROASSIST [zhang2025proactive], EgoSoD [wang2025egosocial], Proactive Agent [luproactive], SocialMind [yang2025socialmind].
6 Enabling Dimensions for Real-World EgoProceAssist
In this section, we formalize two enabling dimensions and introduce their definitions, key methods, datasets, and evaluation protocols, respectively.
6.1 Real-time and Streaming Video Understanding
6.1.1 Definition
Real-time and streaming video understanding is an intelligent analysis paradigm for dynamic, continuous video streams, thereby enabling perception and interpretation of ongoing visual content. By moving beyond the traditional offline setting of processing complete videos in a single pass and instead aligning with the generate-and-process simultaneously nature of real-world video, it has become a key enabling technology for applications such as autonomous driving [zhong2025cooptrack], real-time surveillance, and robotic interaction [yao2025inference].
Rising demand for online, real-time video processing has focused attention on real-time and streaming video understanding and related tasks [de2016online, zhao2022real, wang2021oadtr], such as real-time hand gesture recognition [benitez2021ipn] and 3D pose estimation [luvizon2020multi]. Recent multimodal models further advance real-time interaction across vision, audio, and text [fu2024vita, fu2025vita, xie2024mini]. Aligned with our research, many existing methods already support real-time responses. For example, MiniROAD in PREGO [flaborea2024prego] and TI-PREGO [plini2025ti] are inherently real-time action-detection branches. Likewise, GC [mazzamuto2025gazing] uses gaze prediction to detect errors. Although not always stated, gaze signals are intrinsically streaming and have been experimentally used for online intervention.
Table 7: Performance comparison of representative methods for real-time and streaming video understanding, where metric values are averaged across different tasks within the same dataset. The bold text represents the best performance on this metric. Score (0-5) represents the LLM Judge’s results.
Method Year Accuracy Score (0-5) Coh. F1 Benchmark
VideoLLM-online [chen2024videollm] CVPR 2024 - - - 13.7 ETBench
Dispider [qian2025dispider] CVPR 2025 - - - 31.0
VideoLLM-online [chen2024videollm] CVPR 2023 36.0 - - - StreamingBench
Claude-3.5-Sonnet [bae2024enhancing] - 72.4 - - -
GPT-4o [hurst2024gpt] arXiv 2024 77.1 - - -
Gemini-1.5-Pro [team2024gemini] arXiv 2024 79.0 - - -
Flash-VStream [zhang2024flash] arXiv 2024 23.2 - - -
IXC2.5-OL [zhang2024internlm] arXiv 2024 73.8 - - -
Dispider [qian2025dispider] CVPR 2025 67.6 - - -
Oryx-1.5-7B (StreamBridge) [wang2025streambridge] arXiv 2025 74.8 - - -
LLaVA-OV-7B (StreamBridge) [wang2025streambridge] arXiv 2025 70.9 - - -
Qwen2-VL-7B (StreamBridge) [wang2025streambridge] arXiv 2025 77.0 - - -
Flash-VStream [zhang2024flash] arXiv 2024 33.6 - - - OvO-Bench
Dispider [qian2025dispider] CVPR 2025 41.8 - - -
Oryx-1.5-7B (StreamBridge) [wang2025streambridge] arXiv 2025 71.2 - - -
LLaVA-OV-7B (StreamBridge) [wang2025streambridge] arXiv 2025 69.9 - - -
Qwen2-VL-7B (StreamBridge) [wang2025streambridge] arXiv 2025 71.3 - - -
GPT-4o-mini [hurst2024gpt] arXiv 2024 59.1 3.2 2.0 - StreamBench
VideoLLM-online (FPS: 5) [chen2024videollm] CVPR 2024 56.4 3.1 1.9 -
Flash-VStream (FPS: 1) [zhang2024flash] arXiv 2024 52.1 2.9 2.2 -
StreamChat (FPS: 20) [xiongstreaming] ICLR 2025 63.8 3.4 1.79 -
6.1.2 Methods
As shown in Fig. 7 and Table 7, VideoLLM-online [chen2024videollm] uses streaming EOS prediction to suppress redundant responses and improve inference efficiency, but its single-LLM architecture cannot parallelize perception and reasoning, leading to blocked frame processing during long responses. To address this issue, Dispider [qian2025dispider] decomposes the system into three asynchronous modules, enabling parallel perception and reasoning and alleviating blocking, but at the cost of higher system complexity, greater deployment overhead, and weaker support for short-video offline tasks compared with VideoLLM-online. IXC2.5-OL [zhang2024internlm] performs multimodal streaming over audio and video, extending its applicability to more complex scenarios. However, multimodal processing increases computational and resource costs, and its global-matching memory retrieval performs poorly on fine-grained temporal localization. Flash-VStream [zhang2024flash] addresses this via the STAR memory mechanism, which partitions memory into four complementary sub-modules to compress information and enhance retrieval efficiency. StreamBridge [wang2025streambridge] is a training-free framework that can be directly applied to offline Video-LLMs, using a Round-Decayed Compression strategy to dynamically compress historical content while emphasizing recent information. StreamChat [xiongstreaming] is also training-free and organizes short-term, long-term, and dialogue memory hierarchically to support parallel processing, but its fixed frame-selection threshold risks information loss in highly dynamic scenes. VideoStreaming [qian2024streaming] shares similar limitations. It maintains a constant token length for videos of arbitrary duration, stabilizing memory and inference costs, but cannot adapt the representation length to content complexity, which can lead to redundancy or information loss. MMDuet [wang2024videollm] evaluates each frame’s information value and query relevance in real time to decide whether to interrupt the stream and respond, but it exhibits slow inference and a tendency toward repetitive outputs. To solve it, Video-MoD [wu2024videollm] inserts a LayerExpert module into selected Transformer layers to dynamically select key visual tokens per frame (typically retaining 20%), applying self-attention and feed-forward operations only to these tokens. This reduces redundancy, lowers memory requirements, and improves efficiency.
Table 8: Summarize the representative datasets and benchmarks in real-time and streaming video understanding, Ego means egocentric view and Exo means third-person view.
Dataset Years Duration Videos QA-pairs tasks Domain View
MMDuetIT [wang2024videollm] arXiv 2024 - - - 3 Mixed Ego+Exo
StreamingBench [lin2024streamingbench] arXiv 2024 - 900 4300 18 Mixed Ego+Exo
OVO-Bench [niu2025ovo] CVPR 2025 75h 644 2814 12 Mixed Ego+Exo
SVBench [yangsvbench] ICLR 2025 56.4h 1353 49979 9 Mixed Ego+Exo
Stream-IT [wang2025streambridge] arXiv 2025 - - 120k 8 Mixed Ego+Exo
6.1.3 Datasets
Several related datasets and benchmarks merit particular attention as shown in Table 8. MMDuetIT [wang2024videollm] comprises approximately 109,000 training samples spanning multiple video sources, including Shot2Story [han2023shot2story20k], COIN [tang2019coin], DiDeMo [anne2017localizing], and is further enhanced for real-time streaming video understanding by multi-answer grounded video QA data (Shot2Story-MAGQA-39k) generated with GPT-4o. StreamingBench [lin2024streamingbench] formalizes three core dimensions of streaming video comprehension: real-time visual understanding, holistic source understanding, and contextual understanding, and provides a dataset of 18 tasks, 900 videos, and 4,300 human-annotated QA pairs, spanning eight real-world scenarios such as lifelogging, competitions, education, and more. OVO-Bench [niu2025ovo] systematically evaluates the temporal perception and dynamic reasoning capabilities of VLMs from three key dimensions: backward tracing, real-time visual perception, and forward active responding. It includes 644 unique videos and constructs approximately 2,814 finely grained meta-annotated samples with precise timestamps, covering 12 specific tasks. SVBench [yangsvbench] consists of 1,353 diverse videos sourced from 6 streaming platforms, with an average video length of more than 2 minutes, generating 49,979 high-quality QA-pairs, with an average of 36.94 pairs per video, which is the highest average number of QA-pairs among currently known datasets. Stream-IT [wang2025streambridge] integrates multiple publicly available video annotation datasets (e.g., ActivityNet [caba2015activitynet], Shot2Story [han2023shot2story20k], COIN [tang2019coin]) and incorporates the large-scale synthetic dataset StreamingQA-120K [wang2025streambridge], yielding approximately 1.8 million training samples that span dense video captioning, procedural step recognition, grounded video question answering, and temporal video localization.
6.1.4 Evaluation
Existing works typically evaluate real-time and streaming video understanding via multi-perspective video question answering (VQA), temporal grounding, and event localization. Some also measure models’ processing speed and resource usage on video streams. For question answering or decision-making tasks, accuracy is the dominant metric. Increasingly, the LLM-as-a-Judge paradigm is adopted, using LLMs to score responses. F1, mIoP, mIoU, and TimeDiff assess temporal grounding and event query performance, while Coh. evaluates conversational continuity in multi-turn dialogue. FPS and VRAM are used to quantify inference speed and resource consumption.
6.2 Proactive Interaction in Procedural Contexts
6.2.1 Definition
Proactive interaction is a mode in which an intelligent agent (e.g., a dialogue [zhang2025proactive, deng2023survey] or multimodal assistant [berube2024proactive]) autonomously initiates actions, provides information, or adapts interaction strategies, without explicit user commands, based on continuous perception of context, user state, and task goals, to satisfy explicit or latent user needs. Its timing, content, and form must match the scenario and user expectations to avoid intrusiveness.
With the growing deployment of user-facing assistants [luproactive], research on proactive interaction has emerged across multiple domains [wang2025egosocial]. Proactivity has been extensively studied in social intent analysis [wang2025egosocial, yang2025socialmind, zhang2024proagent] and human–computer interaction [zhang2024proagent] and is gaining attention in work on speech-based voice assistants [berube2024proactive]. Within the three core functions of EgoProceAssist, several methods already support proactive assistance: PREGO [flaborea2024prego] and TI-PREGO [plini2025ti] use predicted future action sets for error detection and can proactively warn users of errors; VINCI [huang2025vinci] performs proactive task planning by suggesting next steps; and HoloAssist dataset [wang2023holoassist] provides explicit labels for proactive interventions. These capabilities make such approaches particularly important for future developments in this field.
6.2.2 Methods
Existing methods for proactive interaction cover many application domains, yet specialized approaches for egocentric vision remain relatively limited. As shown in Fig. 7, SocialMind [yang2025socialmind] captures verbal, non-verbal, and social cues and employs LLM-based reasoning to generate proactive, real-time social suggestions. Tailored to AR hardware, it integrates multimodal signals widely used in first-person vision, thereby aligning closely with egocentric vision research. However, local processing of rich multimodal streams introduces substantial hardware overhead and heightens privacy risks. To reduce this burden, YETI [bandyopadhyay2025yeti] adopts a lightweight, signal-driven paradigm, using video frame structural similarity (SSIM) and object count variation (Alignment Signal) to rapidly decide when to intervene. Also based on LLMs’ reasoning capabilities, ProAgent [zhang2024proagent] proposes a modular Plan-Verify-Control-Memory framework, thereby improving both human-agent collaboration and coordination among multiple agents. Its reliance on predefined task rules and skill libraries, however, restricts generalization to open-ended scenarios, and its memory design offers only limited optimization for long-horizon collaboration. Proactive Agent [luproactive] addresses these memory constraints via a hierarchical architecture that decouples long-term knowledge storage from short-term trajectory memory. EgoSoD [wang2025egosocial] advances proactive interaction by constructing a social cognition graph that integrates audio and visual cues, and by determining intervention timing through a proximity–gaze–speech hierarchical filtering mechanism. PROASSIST [zhang2025proactive] further tackles the severe imbalance between response and non-response frames using Negative Frame Sub-sampling (NFS), and achieves efficient long-video handling through Iterative Progress Summarization (IPS) for dynamic temporal compression and summarization.
Table 9: Summarize the representative datasets and benchmarks in proactive interaction, Ego means egocentric view and Exo means third-person view.
Dataset Years Duration Videos QA-pairs tasks Domain View
ProactiveBench [luproactive] ICLR 2025 - - - 12 Mixed -
ProactiveVideoQA [wang2025proactivevideoqa] arXiv 2025 47.7h 1377 1427 4 Mixed Ego+Exo
STREAMGAZE [lee2025streamgaze] arXiv 2025 66.2h 285 8521 10 Mixed Ego
PROASSIST [zhang2025proactive] arXiv 2025 479h 3934 30135 4 Mixed Ego
EgoSocial [wang2025egosocial] arXiv 2025 4.2h - 13500 - Mixed Ego
6.2.3 Datasets
Besides HoloAssist mentioned above, there are several other important datasets and benchmarks that deserve attention, as shown in Table 9.
ProactiveBench [luproactive] includes 6,790 training events and 233 test events across coding, writing, and daily life, offering large-scale, realistic contexts for training and evaluating the proactive behavior of large language model agents. The authors also built a reward-model subset of 1,760 annotated instances to emulate human acceptance or rejection of task predictions. ProactiveVideoQA [wang2025proactivevideoqa] is the first benchmark focused on evaluating proactive interaction. It combines four tasks: web video QA, egocentric video QA, TV series QA, and video anomaly detection, includes 1,377 videos and 1,427 QA pairs, with about 3.5 turns per video. STREAMGAZE [lee2025streamgaze] defines 10 tasks spanning past, present, and future (including object recognition, gaze-sequence matching, and future action prediction) to comprehensively assess models’ understanding of gaze dynamics and intention prediction, contains 285 videos and 8,521 QA pairs covering real-world interactive scenarios. PROASSIST [zhang2025proactive] constructs dialogue data from the temporal annotations of six public egocentric video datasets, covering a variety of task scenarios such as cooking, object manipulation, and experiments, comprises 30,135 dialogues spanning 479 hours of video, with an average video length of approximately 16 minutes, and provides proactivity and intent labels as well as progress summaries for each assistant turn. EgoSocial [wang2025egosocial] focuses on egocentric social interaction understanding and evaluates multimodal large models’ proactive intervention capabilities in social scenarios. It contains 13,500 video–question pairs annotated with eight categories of social cues, such as speech activity, gaze, and personal space, and it establishes the first quantitative evaluation protocol for social intervention detection.
6.2.4 Evaluation
Proactive model performance is typically evaluated using precision, recall, F1-score, and accuracy to measure the correctness and completeness of predictions. Many studies also tailor metrics to task characteristics. For example, Proactive Agent [luproactive] adds false-alarm rate to capture unnecessary interventions, while EgoSocial [wang2025egosocial] uses intervention timing (accuracy) and macro-F1 to assess precise, non-disruptive interventions across diverse social scenarios. Some approaches further introduce specialized indicators, such as acceptance rate in Proactive Agent [luproactive], or using the LLM-as-a-Judge strategy. GPT-4o, for instance, provides five-point ratings for the correctness, timeliness, and efficiency of proactive behaviors. Human evaluations additionally cover subjective aspects such as conversational naturalness and task alignment, as in SocialMind [yang2025socialmind] and PROASSIST [zhang2025proactive].
Because these methods are evaluated on distinct, domain-specific datasets, their quantitative results are not directly comparable. Accordingly, this section does not include a unified tabular comparison of performance metrics.
@@ 6.2.4 Evaluation
Figure 8: Examples of datasets, EgoPER [lee2024error] and CaptainCook4D [peddi2024captaincook4d] focus on kitchen cooking scenarios, EgoOops [haneji2025egooops] designs complex scenarios such as chemical experiments and building blocks, and EgoproceL [bansal2022my] focuses on pc assembly and disassembly tasks.
@@ 7 Experiments
Table 10: Experimental results in egocentric procedural error detection, the part highlighted in purple is the result of our supplementary experiments, bold text represents the best performance on this metric in each part, and Time is the average inference time. ICL denotes results under an in-context learning setting, and the green upward arrow indicates performance improvement.
Method Years Time Precision EDA AUC F1 Recall Acc Dataset
EgoPED [lee2024error] CVPR 2024 - 56.5 69.8 54.9 - - 69.8 CaptainCook4D
AMNAR [huang2025modeling] CVPR 2025 - 66.8 72.3 60.2 - - 72.3
Just Ask [yang2021just] ICCV 2021 - 35.1 35.1 50.0 51.2 - 35.1
Video-LLaVA [lin2024video] EMNLP 2024 2.55 40.1 60.8 52.2 29.6 23.4 60.8
Video-LLaMA2 [cheng2024videollama] arXiv 2024 4.73 36.5 58.4 50.1 29.8 25.2 58.4
Vinci [huang2025vinci] IMWUT 2025 3.03 35.6 49.7 48.5 - - 49.7
LLaVA-OneVision [lillava] TMLR 2025 4.55 39.8 63.0 53.1 17.0 10.8 63.0
EgoGPT [yang2025egolife] CVPR 2025 3.82 36.0 46.5 49.6 47.0 67.5 46.5
Qwen2.5-Omni-7B [xu2025qwen2] arXiv 2025 4.19 32.3 59.7 53.0 19.1 13.6 59.7
Ming-Lite-Omni [ai2025ming] arXiv 2025 11.5 32.6 51.4 45.7 34.2 36.0 51.4
Video-LLaVA [lin2024video] EMNLP 2024 1.71 14.6 60.0 44.6 15.8 17.1 60.0 EgoOops
Video-LLaVA (ICL) [lin2024video] EMNLP 2024 1.94 20.9 $\uparrow$ 62.5 $\uparrow$ 49.3 $\uparrow$ 23.1 $\uparrow$ 25.7 $\uparrow$ 62.5 $\uparrow$
Video-LLaMA2 [cheng2024videollama] arXiv 2024 3.51 15.4 50.0 42.4 20.0 28.6 50.0
Video-LLaMA2 (ICL) [cheng2024videollama] arXiv 2024 3.59 29.1 $\uparrow$ 57.5 $\uparrow$ 59.4 $\uparrow$ 40.4 $\uparrow$ 65.7 $\uparrow$ 57.5 $\uparrow$
Vinci [huang2025vinci] IMWUT 2025 4.39 19.4 26.9 47.3 30.8 74.3 26.9
Vinci (ICL) [huang2025vinci] IMWUT 2025 4.46 25.5 $\uparrow$ 49.4 $\uparrow$ 58.9 $\uparrow$ 37.2 $\uparrow$ 68.6 49.4 $\uparrow$
LLaVA-OneVision [lillava] TMLR 2025 1.91 26.1 64.4 54.1 29.6 34.3 64.4
LLaVA-OneVision (ICL) [lillava] TMLR 2025 1.89 29.6 $\uparrow$ 71.3 $\uparrow$ 52.6 25.8 22.9 71.3 $\uparrow$
EgoGPT [yang2025egolife] CVPR 2025 3.27 34.8 73.8 55.1 27.6 22.9 73.8
EgoGPT (ICL) [yang2025egolife] CVPR 2025 3.38 30.8 68.8 53.6 32.4 $\uparrow$ 34.3 $\uparrow$ 68.8
Qwen2.5-Omni-7B [xu2025qwen2] arXiv 2025 6.48 10.6 45.6 36.5 13.9 20.0 45.6
Qwen2.5-Omni-7B (ICL) [xu2025qwen2] arXiv 2025 6.50 $\uparrow$ 26.7 $\uparrow$ 48.8 $\uparrow$ 63.4 $\uparrow$ 39.7 $\uparrow$ 77.1 $\uparrow$ 48.8 $\uparrow$
Ming-Lite-Omni [ai2025ming] arXiv 2025 8.03 32.1 59.4 62.8 45.4 77.1 59.4
Ming-Lite-Omni (ICL) [ai2025ming] arXiv 2025 8.15 25.0 41.9 53.0 38.4 82.9 $\uparrow$ 41.9
Video-LLaVA [lin2024video] EMNLP 2024 3.18 13.5 77.5 50.7 14.3 15.3 77.5 EgoPER
Video-LLaMA2 [cheng2024videollama] arXiv 2024 4.71 13.6 75.2 51.6 15.8 18.9 75.2
Vinci [huang2025vinci] IMWUT 2025 4.89 12.0 49.3 47.4 19.3 49.2 49.3
LLaVA-OneVision [lillava] TMLR 2025 7.16 14.0 60.1 54.5 21.3 43.7 60.1
EgoGPT [yang2025egolife] CVPR 2025 7.23 14.4 69.6 51.3 19.4 29.8 69.6
Qwen2.5-Omni-7B [xu2025qwen2] arXiv 2025 9.61 88.2 21.8 47.1 29.0 17.3 21.8
Ming-Lite-Omni [ai2025ming] arXiv 2025 13.2 91.7 18.6 43.1 22.3 12.7 18.6
@@ 7.1 Datasets and Implementation Details
Datasets. To ensure broad scenario coverage, three egocentric datasets named CaptainCook4D [peddi2024captaincook4d], EgoPER [lee2024error], and EgoOops [haneji2025egooops] were employed for egocentric procedural error detection tasks. CaptainCook4D and EgoPER address diverse kitchen activities, while EgoOops targets niche contexts such as chemical experiments and block-building. As these datasets lack predefined test sets, 20% of CaptainCook4D and 30% of EgoOops samples were randomly designated as test sets. For the procedural learning task, EgoProceL [bansal2022my] was used, with PC assembly and disassembly tasks selected to assess the AI assistant’s competence in extracting critical procedural steps.
Implementation Details. Both video question answering models and AI assistants are evaluated using question answering tasks. For egocentric procedural error detection, each model receives 4 to 5 randomly selected questions, with prompts tailored to highlight the error detection objective, thereby maximizing detection performance without fine-tuning. Following the protocol described in [huang2025modeling], keywords from model outputs are parsed to generate confidence scores, which are then used to calculate Precision, EDA, AUC, F1, Recall, and Accuracy. We also measure the average inference time (including sampling) for the first 50 samples. This comprehensive metric set ensures a robust assessment of error detection. Consistent prompt formats and confidence rules are applied across EgoPER and EgoOops, with minor adjustments for CaptainCook4D to accommodate model-specific attributes. To better show how different prompt strategies and context length affect results, we use the EgoOops dataset and test in‑context learning prompts that include definitions of possible error types and brief examples. For egocentric procedural learning, seven models were systematically evaluated on the PC assembly and PC disassembly tasks within EgoproceL, adhering to the experimental protocol established by [mahmood2025procedure]. Performance was quantitatively assessed using the F1 score and Intersection over Union (IoU) metrics, and the results were benchmarked against existing procedural learning methodologies. To address relevant task requirements, tailored prompts were introduced. Given that the models lacked fine-tuning for key step learning and acknowledging the inherent challenge in uniformly evaluating semantically similar key step sequences, the egocentric procedural learning task was appropriately simplified. To mitigate excessive deviation between model predictions and expected outputs, a predefined set of candidate key steps was embedded in the prompts, enabling the models to select the sequence of steps that most closely matched each video segment. In accordance with the specifications of [mahmood2025procedure], and to enhance the objectivity of evaluating procedural learning abilities, the Hungarian matching algorithm was employed to analyze the result. In both experiments, we uniformly sample 16 frames from each step segment as input.